OpenStudy (mattyboyy):

What is the general form of the equation for the given circle centered at O(0, 0)?

jimthompson5910 (jim_thompson5910):

the general equation is (x-h)^2 + (y-k)^2 = r^2

jimthompson5910 (jim_thompson5910):

since the center is (0,0), we know (h,k) = (0,0) so just replace h and k with 0 (x-h)^2 + (y-k)^2 = r^2 (x-0)^2 + (y-0)^2 = r^2 x^2 + y^2 = r^2

jimthompson5910 (jim_thompson5910):

then you replace r with whatever radius you want

OpenStudy (mattyboyy):

@jim_thompson5910 hold on cause I still dont exactly understand

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (mattyboyy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I see

jimthompson5910 (jim_thompson5910):

what is the distance from the center to the edge point?

OpenStudy (mattyboyy):

See thats the thing, I dont remember what my notes said and I cant go back and look at them

jimthompson5910 (jim_thompson5910):

you left them at school or something?

OpenStudy (mattyboyy):

@jim_thompson5910 But I don't want to say its 4,5 cause I just know that is wrong

jimthompson5910 (jim_thompson5910):

hint: 

OpenStudy (mattyboyy):

Ok. So I guess and I got it correct. @jim_thompson5910

OpenStudy (mattyboyy):

@jim_thompson5910 this was the answer x2 + y2 − 41 = 0 but I dont understand honest.

jimthompson5910 (jim_thompson5910):

you use the pythagorean theorem to solve for x

jimthompson5910 (jim_thompson5910):